OpenStudy (anonymous):

determine the equation of the circle which passes through the point (5,4) and is tangent to the y-axis at the point (0,2) ok so i know how an equation of a circle should look Ax^2+Ay^2+Bx+Cy+D=0 and tangent to the y-axis...is what is triggering me off because I would just use the distance formula if it was worded otherwise so....what do i do?

OpenStudy (anonymous):

Tangent to Y axis suggest that the point must be at the half height

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

This show tangent line at various point

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so its all equal to 2^2...but...

OpenStudy (anonymous):

We can plug in two point (5,4) which is given (0,2) which is tangent, (0-a)^2+(2-b)^2 (5-a)^2+(4-b)^2

OpenStudy (anonymous):

now i am confused. if we agree that the center must be at (b,2) than doesn't it follow that \[b=(b-5)^2+(4-2)^2\]? that is the distance between (b,2) and (5, 4) is the same as the distance between (0,2) and (b,2)?

OpenStudy (anonymous):

try as i might i cannot draw a circle tangent to the y-axis at (0,2) that does not have center at (b,2) for some b

OpenStudy (anonymous):

and if i do i get the center at \[\frac{11+\sqrt{5}}{2}\] which is a perfectly good number but seems odd in this context

OpenStudy (anonymous):

i know we have to use the distance formula but..."is tangent to the y-axis at the point (0,2)"

OpenStudy (anonymous):

myininaya found the derivative in terms of y'

OpenStudy (anonymous):

when i draw a circle tangent to the y- axis at (0,2) is surely looks like the center must be at (b,0) i maybe mistaken, but i cannot see it otherwise

OpenStudy (anonymous):

sorry i meant (b,2) not (b,0)

OpenStudy (anonymous):

Okay, these are two equations we get if we plug in points (0,2), (5,4) (0-a)^2+(2-b)^2 (5-a)^2+(4-b)^2 Since center must be (a,2) (0-a)^2+(2-2)^2 (5-a)^2+(4-2)^2 If we set them equal to each other we get a=29/10

OpenStudy (anonymous):

i am willing to admit i am wrong. please show me if it is possible for this not to be so. but if it is true than i think my answer must be correct since the distance between (0,2) and (b,2) is obviously b

OpenStudy (anonymous):

So center is (2.9,2)

OpenStudy (anonymous):

ok you are using a and i used b. i will use a instead.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am so lost...

OpenStudy (anonymous):

If I am right, you found the center (2.9,2)

OpenStudy (anonymous):

ahh ;but i am an idiot because i solved my equation incorrectly. it should be \[a^2=a^2-10a+25+4\]

OpenStudy (anonymous):

and you are right it is a = 2.9

OpenStudy (anonymous):

that's the weird answer

OpenStudy (anonymous):

no that is a normal answer. i solved a quadratic because i forgot that i was using the square of the distance, not the distance. so i wrote b = something rather than b^2=something

OpenStudy (anonymous):

my mistake sorry

OpenStudy (anonymous):

So to recap, here is what we did, 1) Y axis is tangent to circle at mid height, so we were able to find part of the center.(a,2) 2) Using two points, we came up with two equation which we solve for a

OpenStudy (anonymous):

@perplec if you are still confused: 1) the fact that the circle is tangent at (0,2) must mean the center is at (a,2) 2) the fact that it is a circle means the square of the distance between (0,2) and (a,2) which is a^2 is the same as the square of the distance between (a,2) and (5,4) which is \[(a-5)^2+(2-4)^2=a^2-10a+29\]

OpenStudy (anonymous):

setting \[a^2=a^2-10a+29\] gives \[a=\frac{29}{10}=2.9\] and thus the center is \[(2.9,2)\]

OpenStudy (anonymous):

so radius is simply 2.9

OpenStudy (anonymous):

right that is the distance, and the equation is \[(x-2.9)^2+(y-2)^2=2.9^2\]

OpenStudy (anonymous):

so the equation of the circle is? plugging the point (2.9,2) and (5,4) and equating them with the distance formula to find the equation of the circle?

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks so much satellite and imranmeah!!!